The width of a rectangular room is 3 feet less than the length. If the area of the room is 70 square feet, find the length and width of the room.
Answer by TimothyLamb(4379)

w = L - 3
Lw = 70
---
Lw = 70
L(L - 3) = 70
LL - 3L = 70
LL - 3L - 70 = 0
---
the above quadratic equation is in standard form, with a=1, b=-3, and c=-70
---
to solve the quadratic equation, plug this:
1 -3 -70
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
10
-7
---
the negative root doesn't make sense for a length, so use the positive root:
---
L = 10
w = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php